# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ V-A2 Membership Attacks
In the context of LLM-based AI agents, S2MIA could be used to exploit the external knowledge databases that these agents rely on, potentially exposing private data embedded in the database. By targeting the membership privacy of these external databases, attackers could infer which specific data points are being accessed or used in the agent’s responses, compromising user confidentiality. This underscores the importance of developing robust privacy protections for external knowledge sources integrated with LLM-based agents, ensuring that sensitive information remains secure and that agents do not inadvertently expose confidential data during user interaction.
V-B Federated & Multi-Agent System Attacks
V-B1 Contagious Recursive Blocking Attacks
@@
Zhou et al. [56] introduce Contagious Recursive Blocking Attacks (Corba), a novel and highly effective attack method targeting Large Language Model-based Multi-Agent Systems (LLM-MASs). While LLM-MASs have demonstrated impressive capabilities in real-world applications, collaborating to complete complex tasks, their security remains largely underexplored. These systems are designed with safety mechanisms, such as alignment techniques to reject harmful instructions, but they are still vulnerable to targeted disruptions. Corba exploits two key properties: its contagious nature, which allows the attack to propagate across various network topologies, and its recursive property, which depletes computational resources over time. The attack typically involves seemingly benign instructions, making it difficult to detect and mitigate using traditional alignment methods. The authors evaluate Corba on two widely used LLM-MASs, AutoGen and Camel, across various topologies and commercial models. Through extensive experiments in both controlled and open-ended interactive environments, they demonstrate the effectiveness of Corba in disrupting agent interactions within complex topology structures and open-source models. These findings underscore the need for robust security measures to safeguard LLM-MASs against disruptive attacks, particularly those that employ subtle and seemingly harmless instructions.
V-B2 Federated LLM Attack
@@
Han et al. [53] introduce FedSecurity, an end-to-end benchmark designed as a supplementary component of the FedML library for simulating adversarial attacks and defense mechanisms within Federated Learning (FL). FedSecurity simplifies the implementation of fundamental FL procedures, such as training and data loading, allowing users to focus on developing and testing their own attack and defense strategies. The benchmark consists of two key components: FedAttacker, which executes various adversarial attacks during FL training, and FedDefender, which implements defensive mechanisms to mitigate these attacks. FedSecurity offers extensive customization options, making it compatible with a wide range of machine learning models, including Logistic Regression, ResNet, and GAN, as well as FL optimizers like FedAVG, FedOPT, and FedNOVA. The framework enables users to explore the effectiveness of various attacks and defenses across different datasets and models, and it supports flexible configuration via a configuration file and APIs. To demonstrate its utility and adaptability, the authors showcase FedSecurity’s application in the federated training of Large Language Models (LLMs), highlighting its potential for addressing complex applications in FL.
@@
In the scope of AI agents leveraging LLM architectures, FedSecurity can simulate adversarial attacks and defense mechanisms within federated settings, ensuring that AI agents are trained securely and robustly. Using FedAttacker, researchers can evaluate how well multi-agent systems perform under adversarial conditions, while FedDefender can help optimize defense strategies to safeguard agent communications. This benchmark is handy for enhancing the security of decentralized AI systems, ensuring that agents can interact safely while maintaining performance even in the presence of malicious attacks.
V-C Social-Engineering & Data-Extraction Attacks
V-C1 Datastore Leakage Attack
@@
Qi et al. [54] investigate the risk of datastore leakage in Retrieval-Augmented Generation (RAG) Language Models (LMs), focusing on the vulnerability that allows adversaries to extract verbatim text data from the datastore of RAG systems. By leveraging the instruction-following capabilities of LMs, the authors demonstrate that adversaries can easily exploit prompt injection to extract sensitive information from instruction-tuned models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2. The exploitability of this vulnerability increases as the model size scales up, with larger models becoming more susceptible. The paper also explores how RAG setup influences data extractability, revealing that failure to utilize contexts in modern LMs effectively can lead to unexpected instructions causing the model to regurgitate sensitive data. The authors propose that position bias elimination strategies can significantly mitigate this vulnerability. Extending their research to production RAG models, the authors demonstrate an attack that achieves a 100% success rate on 25 randomly selected customized GPT models, extracting text data verbatim from a 77,000-word book at a rate of 41%, and a 1,569,000-word corpus at a rate of 3%, using only 100 queries generated by the models themselves.
@@
Regarding AI agents built on LLM platforms, this datastore leakage vulnerability could be exploited to extract sensitive information from the agents’ memory banks, thereby compromising the confidentiality of the data used by the agents. By manipulating input prompts, adversaries could trigger agents to reveal stored knowledge, including private or proprietary information, without needing direct access to the underlying datastore. This highlights the need for robust security measures, such as advanced context management and query filtering, to prevent LLM agents from inadvertently exposing sensitive information during interactions.
V-C2 Social Engineering attack
@@
Kumarage et al. [57] introduce a novel LLM-agentic framework, SE-VSim, designed to simulate social engineering (SE) attacks in multi-turn, chat-based interactions, particularly in the context of conversational agents, such as chatbots powered by large language models (LLMs). The authors highlight that SE detection in multi-turn conversations is significantly more complex than in single-instance interactions due to the dynamic and evolving nature of these conversations. A key challenge in mitigating SE attacks is understanding the mechanisms through which attackers exploit vulnerabilities, as well as how victims’ personality traits influence their susceptibility to manipulation. To explore this, SE-VSim models victim agents with varying psychological profiles to assess the influence of personality traits on vulnerability to SE attacks. Using a dataset of over 1,000 simulated conversations, the authors simulate attack scenarios where adversaries, posing as recruiters, funding agencies, and journalists, attempt to extract sensitive information. The paper further presents a proof of concept, SE-OmniGuard, a personalized protection mechanism that leverages prior knowledge of the victim’s personality, evaluates attack strategies, and monitors information exchanges during conversations to detect and prevent potential SE attempts.
@@
For AI agents equipped with large language model capabilities, this framework can be used to simulate and analyze how AI agents with varying personality traits respond to SE attacks, enabling the identification of vulnerabilities in agent communication protocols. By understanding how different personality traits influence susceptibility, defensive strategies like SE-OmniGuard could be developed to enhance the resilience of multi-agent systems, ensuring that AI agents can protect sensitive information while interacting with users. This highlights the need for incorporating psychological profiling into the design of secure and trustworthy AI agents, particularly in domains where agents handle sensitive or confidential interactions.
@@
TABLE VI: Comprehensive Security Threat Matrix for MCP [58], A2A[59], ANP[60] & ACP[61] Protocols
@@ VI Vulnerabilities & Attacks for MCP and A2A protocols–based Agent Communications
Autonomous agents rely on standardized communication protocols—most notably MCP and A2A to extend language models with external tools and to coordinate multi‐agent workflows. While these protocols greatly enhance flexibility and interoperability, they also introduce rich new attack surfaces at both the prompt‐processing and orchestration layers. In the subsections that follow, we discuss the primary vulnerability classes, ranging from prompt‐based injections and backdoor mechanisms to discovery spoofing and recursive denial‐of‐service, and illustrate how adversaries can subvert protocol semantics, manipulate context, or abuse authentication to compromise agent communications.
VI-A Model‐Control‐Protocol (MCP)

### v2

@@ V-B2 Membership Attacks
In the context of LLM-based AI agents, S2MIA could be used to exploit the external knowledge databases these agents rely on, potentially exposing private data embedded in them. By targeting the membership privacy of these external databases, attackers could infer which specific data points are being accessed or used in the agent’s responses, compromising user confidentiality. This underscores the importance of developing robust privacy protections for external knowledge sources integrated with LLM-based agents, ensuring that sensitive information remains secure and that agents do not inadvertently expose confidential data during user interaction.
V-C Federated & Multi-Agent System Attacks
V-C1 Contagious Recursive Blocking Attacks
@@
Zhou et al. [zhou2025corba] introduce Contagious Recursive Blocking Attacks (Corba), a novel and highly effective attack method targeting Large Language Model-based Multi-Agent Systems (LLM-MASs). While LLM-MASs have demonstrated impressive capabilities in real-world applications, collaborating to complete complex tasks, their security remains largely underexplored. These systems are designed with safety mechanisms, such as alignment checks to reject harmful instructions, but they remain vulnerable to targeted disruptions. Corba exploits two key properties: its contagious nature, which allows the attack to propagate across various network topologies, and its recursive property, which depletes computational resources over time. The attack typically involves seemingly benign instructions, making it difficult to detect and mitigate using traditional alignment methods. The authors evaluate Corba on two widely used LLM-MASs, AutoGen and Camel, across various topologies and commercial models. Through extensive experiments in both controlled and open-ended interactive environments, they demonstrate the effectiveness of Corba in disrupting agent interactions within complex topology structures and open-source models. These findings underscore the need for robust security measures to safeguard LLM-MASs against disruptive attacks, particularly those that employ subtle and seemingly harmless instructions.
V-C2 Federated LLM Attack
@@
Han et al. [han2024fedsecuritybenchmarkingattacksdefenses] introduce FedSecurity, an end-to-end benchmark designed as a supplementary component of the FedML library to simulate adversarial attacks and defense mechanisms in Federated Learning (FL). FedSecurity simplifies the implementation of fundamental FL procedures, such as training and data loading, allowing users to focus on developing and testing their own attack and defense strategies. The benchmark consists of two key components: FedAttacker, which executes various adversarial attacks during FL training, and FedDefender, which implements defensive mechanisms to mitigate these attacks. FedSecurity offers extensive customization options, making it compatible with a wide range of machine learning models, including Logistic Regression, ResNet, and GANs, as well as FL optimizers such as FedAVG, FedOPT, and FedNOVA. The framework enables users to explore the effectiveness of various attacks and defenses across different datasets and models, and it supports flexible configuration via a configuration file and APIs. To demonstrate its utility and adaptability, the authors showcase FedSecurity’s application in federated training of Large Language Models (LLMs), highlighting its potential to address complex applications in FL.
@@
In the scope of AI agents leveraging LLM architectures, FedSecurity can simulate adversarial attacks and defense mechanisms within federated settings, ensuring that AI agents are trained securely and robustly. Using FedAttacker, researchers can evaluate how well multi-agent systems perform under adversarial conditions, while FedDefender can help optimize defense strategies to safeguard agent communications. This benchmark helps enhance the security of decentralized AI systems, ensuring agents can interact safely while maintaining performance even under malicious attacks.
Attack Category Typical Privilege Feasibility (Low / Medium / High) Primary Layer Impacted
Input Manipulation Attacks (e.g., Prompt Injection, Jailbreaking, Context Hijacking) None to Low (user-level or public API access) High — easily reproducible through open LLM endpoints and exposed toolchains Tool / Agent
Model Compromise Attacks (e.g., Backdoor, Data Poisoning, Memory Injection) Medium to High (training, fine-tuning, or data pipeline access) Medium — requires control over model training or update workflow Host / Model
System & Privacy Attacks (e.g., Membership Inference, Retrieval Poisoning, Side Channels) Medium (runtime or federated environment access) Medium — feasible under shared or multi-agent execution contexts Host / Agent
Protocol Vulnerabilities (e.g., MCP, ACP, ANP, A2A Exploits) Low to Medium (connector, API, or communication access) High — affects integration boundaries and message exchanges between agents Protocol / Agent
TABLE VI: Summary of Attacker Privilege, Feasibility, and Layer Impact Across Major Attack Categories in LLM-Agent Ecosystems. Typical Privilege: Minimum access level required by the adversary (e.g., none, user, or system). Feasibility: Practical likelihood of executing or reproducing the attack in real-world scenarios (Low, Medium, or High). Primary Layer Impacted: Key subsystem affected within the LLM-agent architecture, encompassing the Host, Tool, Agent, or Protocol layer.
V-D Social-Engineering & Data-Extraction Attacks
V-D1 Datastore Leakage Attack
@@
Qi et al. [qi2024follow] investigate the risk of datastore leakage in Retrieval-Augmented Generation (RAG) Language Models (LMs), focusing on the vulnerability that allows adversaries to extract verbatim text data from the datastore of RAG systems. By leveraging LMs’ instruction-following capabilities, the authors demonstrate that adversaries can easily exploit prompt injection to extract sensitive information from instruction-tuned models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2. The exploitability of this vulnerability increases as model size scales up, making larger models more susceptible. The paper also explores how RAG setup influences data extractability, revealing that failing to effectively utilize context in modern LMs can lead to unexpected instructions that cause the model to regurgitate sensitive data. The authors propose that position bias elimination strategies can significantly mitigate this vulnerability. Extending their research to production RAG models, the authors demonstrate an attack that achieves a 100% success rate on 25 randomly selected customized GPT models, extracting text data verbatim from a 77,000-word book at a rate of 41%, and a 1,569,000-word corpus at a rate of 3%, using only 100 queries generated by the models themselves.
@@
Regarding AI agents built on LLM platforms, this datastore leakage vulnerability could be exploited to extract sensitive information from the agents’ memory banks, thereby compromising the confidentiality of the data they use. By manipulating input prompts, adversaries could trigger agents to reveal stored knowledge, including private or proprietary information, without needing direct access to the underlying datastore. This highlights the need for robust security measures, such as advanced context management and query filtering, to prevent LLM agents from inadvertently exposing sensitive information during interactions.
V-D2 Social Engineering attack
@@
Kumarage et al. [kumarage2025personalized] introduce a novel LLM-agentic framework, SE-VSim, designed to simulate social engineering (SE) attacks in multi-turn, chat-based interactions, particularly in the context of conversational agents, such as chatbots powered by large language models (LLMs). The authors highlight that SE detection in multi-turn conversations is significantly more complex than in single-instance interactions, due to the dynamic, evolving nature of these conversations. A key challenge in mitigating SE attacks is understanding the mechanisms through which attackers exploit vulnerabilities, as well as how victims’ personality traits influence their susceptibility to manipulation. To explore this, SE-VSim models victim agents with varying psychological profiles to assess the influence of personality traits on vulnerability to SE attacks. Using a dataset of over 1,000 simulated conversations, the authors simulate attack scenarios where adversaries, posing as recruiters, funding agencies, and journalists, attempt to extract sensitive information. The paper further presents a proof-of-concept, SE-OmniGuard, a personalized protection mechanism that leverages prior knowledge of the victim’s personality, evaluates attack strategies, and monitors information exchanges during conversations to detect and prevent potential SE attempts.
@@
For AI agents equipped with large language model capabilities, this framework can simulate and analyze their responses to SE attacks, enabling the identification of vulnerabilities in communication protocols. By understanding how different personality traits influence susceptibility, defensive strategies like SE-OmniGuard could be developed to enhance the resilience of multi-agent systems, ensuring that AI agents can protect sensitive information while interacting with users. This highlights the need for incorporating psychological profiling into the design of secure and trustworthy AI agents, particularly in domains where agents handle sensitive or confidential interactions.
V-E Attacker Privileges, Feasibility, and Layer Impact Analysis
While the preceding subsections summarized individual research efforts across multiple attack domains, it is equally important to contextualize how these attacks differ in terms of adversarial capability, real-world feasibility, and the specific system layers they impact. This subsection introduces a cross-domain analysis that links the four major categories in our taxonomy—Input Manipulation, Model Compromise, System and Privacy Attacks, and Protocol Vulnerabilities—to their operational requirements and affected components within the LLM-agent ecosystem.
From a privilege perspective, most input manipulation attacks (e.g., prompt injections [perez2022ignore, pedro2023prompt], adaptive indirect injections [zhan2025adaptive], or compositional jailbreaks [jiang2025deceiving]) require no special access to the model or infrastructure. These are typically carried out through public interfaces, API endpoints, or interactive tools exposed to end users. Conversely, model compromise attacks (e.g., backdoors [huang2023composite, zhu2025demonagent], data poisoning [zou2024poisonedrag, alber2025medical]) demand moderate to high privileges, as they exploit training data pipelines or internal model weights. In contrast, system and privacy attacks occupy a middle ground: they often target runtime memory, retrieval pipelines, or federated model states—areas accessible only to semi-trusted agents or insiders [li2024seeing, soleimani2025wiretapping]. Finally, protocol vulnerabilities such as those affecting MCP, ACP, ANP, or A2A involve low to moderate privileges but exhibit high exploitability, as shown in real-world cases such as the Toxic Agent Flow attack or recursive blocking exploits [zhou2025corba].
Feasibility is similarly uneven across categories. Input manipulation and protocol-level exploits are highly feasible [yu2023gptfuzzer, schwartz2025graph], given their reproducibility via public LLM endpoints and the widespread use of loosely validated APIs. Model compromise and system-level attacks, although more sophisticated, are increasingly practical in shared or open-source environments where fine-tuning, memory access, or federated participation is exposed [dong2025practical, han2024fedsecuritybenchmarkingattacksdefenses]. This asymmetry underscores the layered risk landscape: while high-impact attacks may require privileged access, low-privilege attacks remain far more scalable and frequent. Therefore, the affected layers vary as well. Input manipulation attacks primarily compromise the tool and agent layers, influencing behavior through external inputs. Model compromise targets the host or model layers, manipulating training dynamics or parameter space. System and privacy attacks extend into the host and agent layers, exploiting runtime environments or shared retrieval caches. Finally, protocol vulnerabilities operate at the protocol–agent boundary, where data, authentication, or control signals are exchanged between agents and tools [hou2025model, yang2025survey].
Table VI summarizes these contextual dimensions—privilege level, feasibility, and affected layers—providing a practical mapping that complements the detailed literature review. This structured view supports threat modeling, helps prioritize mitigations based on attacker capabilities, and identifies which parts of the LLM-agent stack require urgent security reinforcement. As shown in Table VI, attacks that require minimal privileges—such as prompt injections or protocol exploits—remain the most prevalent and replicable in practice, posing a disproportionate security threat [perez2022ignore, zhan2025adaptive]. In contrast, deeper model-compromise or data-poisoning attacks, while harder to execute, carry severe long-term consequences by undermining the trustworthiness and alignment of the underlying LLM [huang2023composite, zhu2025demonagent]. This analysis reinforces the need for multi-layered defense strategies that span access control, input validation, model integrity verification, and secure protocol design [hou2025model, wang2025comprehensive].
@@ V-E Attacker Privileges, Feasibility, and Layer Impact Analysis
TABLE VII: Comprehensive Security Threat Matrix for MCP [modelcontextprotocol2025], A2A[googleblog2025], ANP[agentnetworkprotocol2024] & ACP[agentcommunicationprotocol2024] Protocols
@@ VI Vulnerabilities & Attacks for MCP and A2A protocols–based Agent Communications
Autonomous agents rely on standardized communication protocols—most notably MCP and A2A — to extend language models with external tools and coordinate multi-agent workflows. In this work, the term protocol vulnerabilities refers specifically to weaknesses in agent coordination and communication standards rather than general network protocols. These include the Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent Network Protocol (ANP), and Agent-to-Agent (A2A) protocol, which govern structured interactions between LLM-powered agents and external tools or services. Such vulnerabilities arise from misconfigured authentication, context manipulation, and insecure orchestration flows, distinguishing them from traditional transport-layer issues (e.g., HTTP or gRPC exploits).
@@
While these protocols greatly enhance flexibility and interoperability, they also introduce rich new attack surfaces at both the prompt‐processing and orchestration layers. In the subsections that follow, we discuss the primary vulnerability classes, ranging from prompt‐based injections and backdoor mechanisms to discovery spoofing and recursive denial‐of‐service, and illustrate how adversaries can subvert protocol semantics, manipulate context, or abuse authentication to compromise agent communications.
VI-A Formal Definition of the Protocol‐Level Threat Model
To rigorously capture the semantics of communication-layer vulnerabilities among autonomous agents, we formalize the Protocol‐Level Threat Model, which encompasses the Model Context Protocol (MCP), Agent Network Protocol (ANP), Agent Communication Protocol (ACP), and Agent-to-Agent (A2A) interactions. These protocols introduce structured message exchanges between language‐model‐driven agents and their tools, databases, or peers. Vulnerabilities at this level arise not from model weights but from insecure message orchestration, authentication misconfigurations, or context propagation flaws across agents. The following formulation defines how such attacks can be represented mathematically and classified across the orchestration hierarchy.
VI-A1 System Model
Let $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$ denote a set of autonomous agents, each equipped with an internal language model $M_{\theta_{i}}$ and a communication interface defined by a protocol $\mathcal{P}_{i}=(\mathcal{M}_{i},\mathcal{K}_{i},\mathcal{R}_{i})$, where:
  • $\mathcal{M}_{i}$ is the message schema or API specification (e.g., MCP request/response, JSON-RPC task object);
  • $\mathcal{K}_{i}$ is the context store (memory, embeddings, or retrieved results);
  • $\mathcal{R}_{i}$ is the rule set governing message sequencing, authentication, and verification.
Each agent transmits or receives structured messages:
$m_{ij}=\langle\text{id},\text{src},\text{dst},\text{type},\text{payload},\sigma\rangle,$ (37)
where $\sigma$ is the authentication or integrity signature. A valid exchange satisfies:
$\text{Verify}_{\mathcal{R}_{j}}(m_{ij},\sigma)=1.$ (38)
The agent output at time $t$ is thus a function of the model state, received messages, and retrieved context:
$y_{i}^{t}=M_{\theta_{i}}(x_{i}^{t},\mathcal{K}_{i}^{t},\mathcal{M}_{i}^{t}).$ (39)
VI-A2 Adversary Model
An adversary $\mathcal{A}^{*}$ can manipulate or intercept protocol messages, impersonate an agent, or alter context values. Formally, the adversary introduces a transformation $g_{p}$ over the message space:
$g_{p}:m_{ij}\mapsto m^{\prime}_{ij}=(\text{id}^{\prime},\text{src}^{\prime},\text{dst}^{\prime},\text{type}^{\prime},\text{payload}^{\prime},\sigma^{\prime}),$ (40)
subject to one or more of the following goals:
Integrity violation: $\quad m^{\prime}_{ij}\neq m_{ij}\wedge\text{Verify}_{\mathcal{R}_{j}}(m^{\prime}_{ij},\sigma^{\prime})=1,$ (41)
Impersonation: $\quad\text{src}^{\prime}\neq\text{src},$ (42)
Context tampering: $\quad\mathcal{K}_{i}^{\prime}=\mathcal{K}_{i}\oplus\delta_{k},$ (43)
Denial of service: $\quad\exists t:\;|\{m_{ij}^{t}\}|>\tau_{\text{rate}},$ (44)
where $\tau_{\text{rate}}$ is the resource threshold. The attack succeeds if the protocol-level policy $\pi_{\text{proto}}(m^{\prime}_{ij})=0$, i.e., the modified exchange violates communication or access control constraints.
VI-A3 Authentication & Credential Attacks
Authentication and token‐theft attacks can be expressed as violations of signature consistency:
$\sigma^{\prime}\neq\text{Sign}_{\text{auth}}(\text{src}^{\prime},\text{payload}^{\prime})\quad\text{and}\quad\text{Verify}_{\mathcal{R}_{j}}(m^{\prime}_{ij},\sigma^{\prime})=1.$ (45)
This occurs when replayed, stolen, or confused‐deputy tokens remain cryptographically valid due to misconfigured verification or missing nonce enforcement in $\mathcal{R}_{j}$.
VI-A4 Cross‐Agent Context Manipulation
Context manipulation attacks exploit unvalidated state propagation between agents. Given a context aggregation function:
$\mathcal{K}_{i}^{t+1}=f(\mathcal{K}_{i}^{t},\text{payload}(m_{ji})),$ (46)
the adversary injects $\delta_{k}$ such that:
$\mathcal{K}_{i}^{t+1}=f(\mathcal{K}_{i}^{t}\oplus\delta_{k},\text{payload}(m_{ji}^{\prime})),$ (47)
resulting in erroneous or malicious context influencing downstream decisions. This mechanism underlies cross-agent prompt injection, discovery spoofing, and replay-based corruption.
VI-A5 Recursive & Contagious Workflow Exploits
Recursive DoS and contagious blocking occur when agents recursively generate tasks:
$T_{i}^{t+1}=g(T_{i}^{t})=\text{Spawn}(A_{j},T_{i}^{t}),$ (48)
creating unbounded task trees $|T|\to\infty$. If resource consumption exceeds $\tau_{\text{cap}}$, the orchestrator halts or crashes, leading to distributed denial of service across the A2A network.
VI-A6 Protocol‐Level Threat Composition
The complete protocol threat model can be expressed as:
$\mathcal{G}_{\text{proto}}=\mathcal{G}_{\text{auth}}\cup\mathcal{G}_{\text{ctx}}\cup\mathcal{G}_{\text{dos}},$ (49)
representing the union of authentication/credential attacks ($\mathcal{G}_{\text{auth}}$), context‐manipulation exploits ($\mathcal{G}_{\text{ctx}}$), and recursive/DoS vectors ($\mathcal{G}_{\text{dos}}$). Each submodel corresponds to a distinct layer of the agent‐communication pipeline, mapping MCP to vertical integration (tool invocation) and A2A to horizontal collaboration (multi‐agent orchestration). Together, they define the uppermost layer of the end‐to‐end LLM‐Agent Security Stack.
VI-B Model‐Control‐Protocol (MCP)